OpenStudy (anonymous):

Colin needs to find negative factors of 336 that have a difference of 8 units. Which equation could Collin use to find the smaller negative factor? x2 + 8x – 336 = 0 x2 – 8x + 336 = 0 x2 – 4x – 328 = 0 x2 + 4x + 328 = 0

OpenStudy (anonymous):

@Nnesha @wio @robtobey @hartnn @perl

OpenStudy (perl):

first produce a system of equations for the direction

OpenStudy (anonymous):

How

OpenStudy (perl):

find negative factors of 336 that have a difference of 8 units. Assume x is greater than y x*y = 336 x - y = 8 solve this system

OpenStudy (anonymous):

What is x and y

OpenStudy (perl):

they are unknowns

OpenStudy (perl):

You could label them a, b if you want

OpenStudy (anonymous):

So thats the equation

OpenStudy (perl):

it might be able to answer this easier if you look at the given choices

OpenStudy (perl):

`x^2 + bx + c = 0` to factor this you have to find two factors of `c` that add up to `b`

OpenStudy (anonymous):

So either A or B

OpenStudy (perl):

right

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@perl ^

OpenStudy (perl):

yes that sounds correct

OpenStudy (perl):

since two negatives make a positive when you multiply

OpenStudy (anonymous):

Right

OpenStudy (perl):

we want two negative factors of 336 , and (-)(-) = +

OpenStudy (anonymous):

Right